Construct and explain a histogram regarding nationwide officer involved shootings resulting in fatalities using the following data 2015 = 995, 2016 = 963, 2017 = 987, and 2018 = 998

  1. frequency distribution table for the data selected (includes cumulative frequency, relative frequency, and cumulative relative frequency)
  2. histogram graph based on the frequency distribution
  3. summary paragraph discussing observations and conclusions
Answer #1

Solution:-

Year Frequency Relative Frequency Cumulative Frequency Cumulative relative Frequency
2015 995 995/3943=
0.252346
995 0.252346
2016 963 963/3943=
0.24423
995+963=1958 0.252346+0.24423=0.496576
2017 987 987/3943=
0.250317
1958+987=2945 0.496576+0.250317=0.746893
2018 998 998/3943=
0.253107
2945+998=3943 0.746893+0.253107=1
Total
3943
1

Histogram:-

1250 2015 Year: 99!5 998 987 1000 963 矜. 500 250 2015 2017 2018 Year

from the graph we can conclude that the nationwide officers involved shootings resulting in fatalities distribution are distributed Uniformly.